Frank is lending $1,000 to Sarah for two years. Frank and Sarah agree that Frank should earn a real return of 2 percent per year.

Instructions: Enter your responses as whole numbers.

a. The CPI (times 100) is 100 at the time that Frank makes the loan. It is expected to be 104 in one year and 108.2 in two years. What nominal rate of interest should Frank charge Sarah?

The nominal rate of interest charged should be  %.

b. Suppose Frank and Sarah are unsure about what the CPI will be in two years. How should Frank index Sarah's annual repayments to ensure that he gets an annual 2 percent rate of return?

Frank should charge Sarah  %   (Click to select)   equal to   more than   less than  the inflation rate.

You have just graduated from college. You, your older sister, your mother and your father are comparing the starting salaries in your first jobs. The year you each started your first job, the CPI in that year and your staring salary are given below. Which of you started your first job with the highest real income?

Instructions: Enter your responses rounded to two decimal places.

Year Family member CPI Starting Salary Real Income
1980 Your father 0.82 $13,400 $  
1985 Your mother 1.08 $16,900 $  
2010 Your older sister 2.18 $33,200 $  
2017 You 2.44 $16,900 $  


The family member that started with the highest real income was   (Click to select)   you   your older sister   your father   your mother  .

The family member that started with the lowest real income was   (Click to select)   your father   you   your mother   your older sister  .
Answer #1

1) Nominal interest rate = real rate+inflation

a) Inflation in year 1 = 104-100/100*100 = 4%

Frank should charge a nominal interest rate of 4+2 =6%

b) Frank should charge Sarah 2% more than inflation rate.

As per Chegg guidelines, the first question is answered